The electrical resistance R ohm for a pure metal wire is related to its temperature T (in c^0) by the formula R=R_o (1+aT) for positive constant a and R_o.

For what temperature is R=R_o.

Assuming that the resistance is 0 if T=-270 C^o, find a.

Silver wire has a resistance of 1.25 ohms at 0c^o. At what temperature is the resistance 2 ohms?


1
Expert's answer
2020-12-01T02:08:28-0500

Consider the function for resistance

"R = R_0(1 + \u03b1T)" ohm

Temperature when "R = R_0"

"R_0 = R_0(1 + \u03b1T)"

0= 1 + αT

"T = \frac{-1}{α}"

When R = 0 and T = -273 ºC:

"R_0(1 + \u03b1(-273)) = 0"

1 + α(-273) = 0

"α = \frac{-1}{-273} = \frac{1}{273}"

When R = 0 and T = -273 ºC, the constant "α = \frac{1}{273}"

Now when R = 1.25 and T = 0 ºC:

"R_0(1 + \u03b1(0)) = 1.25"

"R_0(1 + 0) = 1.25"

"R_0 = 1.25"

Now if R = 2, we get:

(1.25)(1 + αT)) = 2

"(1 + \frac{1}{273}T) = 1.6"

"T = 0.6 \times 273 = 163.8\;ºC"

When R = 2 the temperature is T = 163.8 ºC